OpenStudy (anonymous):

solving linear equation -7/8=-5/6z

geerky42 (geerky42):

Divide both sides by -5/6 to isolate z.

OpenStudy (anonymous):

i tried but i got way different answer then book

geerky42 (geerky42):

Can you show me your work?

OpenStudy (anonymous):

-0.875*0.833=z 0.729=z

geerky42 (geerky42):

Let me show you. \[-\dfrac{7}{8} = -\dfrac{5}{6}z\]Divide both sides by \(-\dfrac{5}{6}\). \[\dfrac{-\dfrac{7}{8}}{-\dfrac{5}{6}} = \dfrac{-\dfrac{5}{6}z}{-\dfrac{5}{6}}\] \[\dfrac{-\dfrac{7}{8}}{-\dfrac{5}{6}} = z\] Is this clear?

OpenStudy (anonymous):

but how the answer is 21/20

OpenStudy (anonymous):

now i got it -7/8*6/5=z z=42/40 or 21/20

geerky42 (geerky42):

Glad you got it.

OpenStudy (anonymous):

thanks greeky42

geerky42 (geerky42):

You're welcome.